Question Help Apartice with unit charge (q 1)enter
Answer #1

Given q=1 B= i + j

The magnitude of force is given as F=qvB

= 1 x 19k x i + j = 19k(i +j)

=19 i + 19 j N

F = sqrt (192 + 192)

=26.87 N

According to the right-hand rule

When the magnetic force relationship is applied to a current-carrying wire,

This force will be perpendicular to Magnetic field.